OpenStudy (anonymous):

Will medal Julian fully simplifies this polynomial and then writes it in standard form. 4x2y2 – 2y4 – 8xy3 + 9x3y + 6y4 – 2xy3 – 3x4 + x2y2 If Julian wrote the last term as –3x4, which must be the first term of his polynomial in standard form?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Help pls

OpenStudy (123ab456c):

Answer: =−3x4+9x3y+5x2y2−10xy3+4y4

OpenStudy (123ab456c):

EASY!!

OpenStudy (anonymous):

but what is the first term

OpenStudy (123ab456c):

do you get it or should i show my work

OpenStudy (123ab456c):

Let's simplify step-by-step. 4x2y2−2y4−8xy3+9x3y+6y4−2xy3−3x4+x2y2 =4x2y2+−2y4+−8xy3+9x3y+6y4+−2xy3+−3x4+x2y2 Combine Like Terms: =4x2y2+−2y4+−8xy3+9x3y+6y4+−2xy3+−3x4+x2y2 =(−3x4)+(9x3y)+(4x2y2+x2y2)+(−8xy3+−2xy3)+(−2y4+6y4) =−3x4+9x3y+5x2y2+−10xy3+4y4

OpenStudy (anonymous):

ohhhh thank u

OpenStudy (123ab456c):

can you fan me and medal me

OpenStudy (anonymous):

sure

OpenStudy (123ab456c):

thank you hon!!

OpenStudy (123ab456c):

did u fan?